### v1

@@ 6 Related Work
Scheduling layer components within Large Language Models (LLMs) has become a critical area of research due to the increasing demands of improving the efficiency of these models Ye et al. (2024) Li et al. (2024). Effective scheduling can improve end-to-end network performance when components underutilize GPU resources, if they exhibit distinct resource utilization patterns.
@@
Several studies have explored different aspects of LLM scheduling. For instance, Splitwise Patel et al. (2024) proposes a technique to separate the prefill and decoding phases of LLM inference onto different machines because of their unique characteristics. Similarly, Muxserve Duan et al. introduces a spatial-temporal multiplexing system that can flexibly colocate or separate these phases to maximize runtime efficiency.
@@
A popular scheduling approach involves overlapping computation-intensive components, such as matrix multiplications (matmuls) that utilize floating-point (FP) units, with inter-GPU communication tasks. The large sizes of LLM networks often require the workload to be divided executed in parallel on multiple GPUs, where communication layers are required. Several forms of parallelism have been proposed to enhance efficient model training and serving, including Data Parallelism Li et al. (2020), Tensor Parallelism Shoeybi et al. (2019), Expert Parallelism Rajbhandari et al. (2022)NVIDIA (g), and Sequence Parallelism Li et al. (2021). Since there is dependency between data on different GPUs, parallelism involve essential communication layers that transfer large amount of data between each GPU. Given the substantial communication overhead and distinct resource requirements between the data-transfer-heavy communication layers and the computation-heavy GEMM layers, it is ideal to overlap them to improve runtime. Best practices suggest splitting these components into finer-grained chunks and pipelining them for efficient overlapping Pati et al. (2024)NVIDIA (b). This method is often enhanced by fine-grained kernel fusion techniques to further optimize performance Chang et al. (2024)Punniyamurthy et al. (2023).
@@
In addition, several theoretical frameworks have been developed to model and analyze LLM workload performance by pinpointing hardware bottlenecks. Examples include roofline models Yuan et al. (2024), large-scale simulation framework Agrawal et al. (2024b)Agrawal et al. (2024a), and light-weight performance modeling approaches Zhang et al. (2024). Our work uses a similar approach of analyzing LLM training performance based on hardware constraints, and extends these methodologies by providing detailed insights into the effectiveness of layer overlapping strategies.
7 Conclusion
@@
This paper proposed a strategic overlapping of RNG with GEMM layers to improve end-to-end LLM training efficiency. By decoupling RNG from the Dropout process and running it in parallel with the computationally intense GEMM operations, we effectively reduce the latency impact typically associated with RNG if fused with the Attention layer. Our approach optimizes the use of hardware resources by exploiting the distinct hardware demands of RNG and GEMM, and achieves a notable improvement in end-to-end training performance, with speedups ranging from 1.06x to 1.23x within a Transformer Block across various LLM architectures.
@@
We also develop a fine-grained theoretical performance model, validated with silicon results, to provide deeper insights into overlapping different kernels in the LLM network and highlight potential areas for further improvements. The principles established here can extend to optimizing other network layers, offering a generalized strategy to analyze the implications and enhance the performance of future LLM systems. As LLMs continue to scale and the demand for computational efficiency grows, the theoretical model can serve as a valuable framework for evaluating future overlapping strategies maximize resource utilization and minimize training time.

### v2

@@ 8. Related Work
Scheduling layer components within Large Language Models (LLMs) has become a critical area of research due to the increasing demands of improving the efficiency of these models (Ye et al., 2024) (Li et al., 2024). Effective scheduling can improve end-to-end network performance when components underutilize GPU resources, if they have distinct resource utilization patterns.
@@
Several studies have explored different aspects of LLM scheduling. For instance, Splitwise (Patel et al., 2024) proposes a technique to separate the prefill and decoding phases of LLM inference onto different machines because of their unique characteristics. Similarly, Muxserve (Duan et al., [n. d.]) introduces a spatial-temporal multiplexing system that can flexibly co-locate or separate these phases to maximize runtime efficiency.
@@
A popular scheduling approach involves overlapping computation intensive components, such as matrix multiplications (matmuls) that utilize floating-point (FP) units, with inter-GPU communication tasks. The large sizes of LLM networks often require the workload to be divided and executed in parallel on multiple GPUs, where communication layers are required. Several forms of parallelism have been proposed to enhance efficient model training and serving, including Data Parallelism (Li et al., 2020), Tensor Parallelism (Shoeybi et al., 2019), Expert Parallelism (Rajbhandari et al., 2022)(NVIDIA, 2021), and Sequence Parallelism (Li et al., 2021). Since there is dependency between data on different GPUs, parallelism involves essential communication layers that transfer large amount of data between each GPU. Given the substantial communication overhead and distinct resource requirements between the data-transfer-heavy communication layers and the computation-heavy GEMM layers, it is ideal to overlap them to improve runtime. Best practices suggest splitting these components into finer-grained chunks and pipelining them for efficient overlapping (Pati et al., 2024)(NVIDIA, 2025a). This method is often enhanced by fine-grained kernel fusion techniques to further optimize performance (Chang et al., 2024)(Punniyamurthy et al., 2023).
@@
In addition, several theoretical frameworks have been developed to model and analyze LLM workload performance by pinpointing hardware bottlenecks. Examples include roofline models (Yuan et al., 2024), large-scale simulation framework (Agrawal et al., 2024b)(Agrawal et al., 2024a), and light-weight performance modeling approaches (Zhang et al., 2024). Our work uses a similar approach of analyzing LLM training performance based on hardware constraints, and extends these methodologies by providing detailed insights into finer-grained architecture details to evaluate the effectiveness of layer fusion and overlapping strategies.
9. Conclusion
@@
This paper proposed a strategic overlapping of RNG with GEMM layers to improve end-to-end LLM training efficiency. By decoupling RNG from the Dropout process and running it in parallel with the computationally intense GEMM operations, we effectively reduce the latency impact typically associated with RNG if fused with the Attention layer. Our approach optimizes the use of hardware resources by exploiting the distinct hardware demands of RNG and GEMM, and achieves a notable improvement in end-to-end training performance, with 1.22x speedup within a Transformer Block for Llama3 and similar improvements across various LLM architectures.
@@
We develop a fine-grained analytical performance model, validated with silicon results, to provide deeper insights into overlapping different kernels in the LLM network and highlight potential areas for further improvements. The principles established here can extend to optimizing other network layers, offering a generalized strategy to analyze the implications and enhance the performance of future LLM systems. As LLMs continue to scale and the demand for computational efficiency grows, the theoretical model can serve as a valuable framework for evaluating future overlapping strategies maximize resource utilization and minimize training time.